ITF found locked in LOW_NOISE_3 with SCIENCE mode (Autoscience ON); BNS Range ~54 Mpc. Violin modes excited after last relock due to earthquake (DMS plot in fig.1). Unlocks (times in UTC):
02:38 - from LN3 due to PR_TX missing data? (pdf 1). 02:58 - from CARM_NULL_1F with a long series of IMC unlocks for the next handful of minutes. 03:40 - from ACQUIRE_LOW_NOISE_3. 03:44 - from LOCKING_ARMS_IR.
ITF relocked at 04:31UTC, back in SCIENCE at 04:34UTC. Unlocked again at 04:55UTC due to diverging SSFS correction(? pdf 2) relock in progress.



